OpenStudy (anonymous):

5√(- 1/32) I come out with 5i√(2)/8 what am I doing wrong

OpenStudy (anonymous):

I think you should forget about i unless there are square roots (i^2 =1) or i is given in the question.

OpenStudy (anonymous):

You have: \[5\sqrt{\frac{-1}{32}}=5i \sqrt{\frac{1}{32}}=5i \sqrt{\frac{1}{4*8}}=\frac{5i}{2}\sqrt{\frac{1}{8}}\]

OpenStudy (anonymous):

\[\frac{5i}{2}\sqrt{\frac{1}{4*2}}=\frac{5i}{2}*\frac{1}{2}\sqrt{\frac{1}{2}}=\frac{5i}{4}\sqrt{\frac{1}{2}}\]

OpenStudy (anonymous):

Then you can rewrite that as \[\frac{5i}{4*\sqrt{2}}*\frac{\sqrt{2}}{\sqrt{2}}=\frac{5\sqrt{2}i}{8}\]

OpenStudy (anonymous):

Does that make sense?

OpenStudy (anonymous):

Why isn't the fifth root of -1 just -1? (so -1/2)?

OpenStudy (anonymous):

Is it 5th root or 5*sqrt(whatever)?

OpenStudy (anonymous):

where's the fraction symbols

OpenStudy (anonymous):

\[\sqrt[5]{-1/32}\]

OpenStudy (anonymous):

-1/2 (fifth root of -1 is -1 and fifth root of 32 is 2.

OpenStudy (anonymous):

He is correct. I was assuming 5*sqrt(whatever)